e alulátión of equivalent units Use the following information to answer questions 13, 14 and 15: Shitfon Electronics manufactures music players. Its costing system uses two cost direct materials and conversion costs. Each product must pass through the Assembly Department, the Programming department, and the Testing Department. added at the beginning of the production process. Con throughout production. Shiff information is available for the month of March for the Assembly department. Direct materials are version costs are allocated evenly on Electronics uses weighted-average costing. The following ts (40% complete as to conversion costs) Work in process, beginning inventory Units started during March Work in process, ending inventory 890 units 100 units (50% complete as to conversion costs) ら 3,0 The cost details for the month of March are as follows: Work in process, beginning inventory 101. CC Direct materiats Conversion costs 346 $366,000 Direct-meterials costs added during March $703,500 Conversion eosts added during March S1,128,500 요) 13. What are the equivalent units for conversion costs for March? e d 44230 units 1.5 130 units 180 unitsX D) 1,330 units 14.What is the conversion cost per equivalent unit in March? (A S1,266.53 $1,215.04 C) $917.48 D) $956.36 15. What amount of conversion costs is assigned to the units completed from the Assembly department for March? 1,797,178.9 旷$1,127,211.7 $63,326.50 D)$1,431,178.90-
Answer #1

13) Equivalent unit of conversion = 1130+(100*50%) = 1180 units

so answer is c) 1180 units

14) Cost per equivalent unit = (366000+1128500)/1180 = 1266.53 per equivalent unit

so answer is a) $1266.53

15) Conversion cost of units completed = 1266.53*1130 = 1431178.90

so answer is d) $1431178.90